# Viscosity of polymer melts using non-affine theory based on vibrational modes

### v1

@@ II.2 Green-Kubo method for viscosity calculation
The Green-Kubo (GK) method [33, 34] is the most widely used technique to calculate transport coefficients, such as heat conductivity [76], shear viscosity, and bulk viscosity, relating the time correlation functions to the corresponding fluxes or tensors in thermal equilibrium. To calculate the shear viscosity using this method, at first the autocorrelation function of stress $\sigma_{\alpha\beta}$ (with $\alpha,\beta=x,y,z$) is calculated as a function of time with the equilibrium trajectory of the system. The viscosity is then obtained by integrating the stress autocorrelation function over time, and the expression is given by [77]
@@ II.3 Viscosity via non-equilibrium molecular dynamics
Another very popular method to estimate viscosity is via non-equilibrium molecular dynamics (NEMD). In this approach, the system is subjected to shear deformation at a rate $\dot{\gamma}$, and the resulting shear stress is measured as the deformation progresses. Initially, the response is linear, followed by a nonlinear regime, until the system reaches a steady state where the shear stress fluctuates around a mean value. If the system is sheared along the $xy$-plane, the stress response is given by $\sigma_{xy}$, and the steady-state average shear stress for a given shear rate is denoted as $\langle\sigma_{xy}\rangle_{\rm S}$. The shear viscosity $\eta$ is then defined as: $\eta=\langle\sigma_{xy}\rangle_{\rm S}/\dot{\gamma}$.
When viscosity is plotted as a function of shear rate, its behavior depends on the nature of the fluid. If viscosity decreases with increasing shear rate, the material exhibits shear-thinning behavior [68], whereas an increase in viscosity with shear rate indicates shear-thickening behavior [78]. A constant viscosity across shear rates characterizes Newtonian behavior. For most of the complex fluids, shear-thinning/shear-thickening vanishes at sufficiently small shear-rates and viscosity becomes a constant independent of shear-rate. Such a viscosity is called zero-shear viscosity given by,
@@ III Results and Discussion
We applied the framework discussed in Section II.1, utilizing the microscopic formula Eq. (6) to establish the relationship between viscosity and normal mode analysis. The vibrational density of states is plotted in Fig. 1(a) for three different temperatures. Since our system is at a finite temperature, the vDOS has a significant number of imaginary modes (conventionally shown on the negative branch of $\omega$) along with the real modes[61, 62, 64]. The imaginary modes appear due to the presence of saddles in the energy landscape. As temperature decreases, the number of imaginary modes also decreases, whereas the number of real modes increases. There are two peaks on the positive side of the frequency axis, largest peak occurring at a smaller frequency (close to $\omega=10$) is attributed to the vibration arising from the LJ interactions between beads, while the smaller peak is slightly broader and occurs at higher frequency (close to $\omega=50$), primarily attributed to the vibrations corresponding to non-breakable FENE bonds [79, 67]. As expected, $g(\omega)$ exhibits a linear behavior at low frequencies, consistent with both theoretical predictions [69] and experimental observations [70, 71].
@@
Figure 1: (a) Vibrational density of states $g(\omega)$ and (b) affine force field correlator $\Gamma(\omega)$ as a function of normal mode frequency $(\omega)$ at different temperatures (marked). The numerical calculation is done for the KG polymer system described in the text.
@@
In Fig. 1(b), the affine force field correlator $\Gamma(\omega)$ is plotted as a function of the normal mode frequency $\omega$ for three different temperatures. We observe that $\Gamma(\omega)$ starts increasing rapidly for $\omega>20$ and exhibits a temperature-dependent peak near $\omega\simeq 70$. The peak intensity decreases as the temperature decreases and shifts towards lower frequencies. The decay of the correlator function is also temperature-dependent, faster decay is observed at smaller temperatures. At low frequencies, $\Gamma(\omega)$ is expected to follow the analytical behavior[46, 80] $\Gamma\sim\omega^{2}$.
Figure 2: Dependence of the (a) real ($G^{\prime}$, storage modulus) and (b) imaginary ($G^{\prime\prime}$, loss modulus) parts of the complex viscoelastic modulus $(G^{*})$ on external deformation frequency $\Omega$ at different temperatures (marked).
@@
Figure 3: (a) Stress autocorrelation function at different temperatures (marked). Inset shows the tail part of the correlation function. (b) Plot of Green-Kubo viscosity as a function of time at different temperatures (marked).
@@
Using the information of $g(\omega)$ and $\Gamma(\omega)$ we calculate the viscoelastic modulus as a function of external frequency $\Omega$ in the liquid and supercooled regions. In Fig. 2 we plot real and imaginary parts, $G^{\prime}$ and $G^{\prime\prime}$, which represent the storage and loss moduli respectively, as a function of external frequency $\Omega$ at different temperatures. $G^{\prime}$ exhibits specific features, such as, there exists two different resonating frequencies where the response is significantly high with respect to the response at nearby frequencies. The resonating peak at a lower frequency occurs at a frequency where $\Omega$ matches with the vibrational eigenfrequency of LJ interaction. The other peak is higher than the previous peak and occurs at a higher frequency, close to the vibrations corresponding to FENE bonds[50, 67]. Since these bonds are non-breakable, the mechanical response at $\Omega$ corresponding to FENE bonds is strongest. Below the resonance frequency, $G^{\prime}$ decreases and converges to a temperature-dependent finite value in the low-frequency regime, where $G^{\prime}$ is proportional to $\Omega^{0}$, consistent with the Zener standard linear solid model of viscoelasticity [41, 26]. Also, at $\Omega$ greater than resonating frequencies, there exists a temperature-dependent plateau, representing the high-frequency modulus dominated by affine displacements.
@@
In Fig. 2(b), we plot $G^{\prime\prime}$ as a function of $\Omega$ at three temperatures. We observe that the value of $G^{\prime\prime}$ becomes very small at both low and large frequencies. It exhibits a broad peak at $\Omega\sim 70$, and below this frequency $G^{\prime\prime}$ decreases, forming a plateau having a wider width at smaller temperatures due to slower decay in the low-frequency regime. Subsequently, $G^{\prime\prime}$ decreases linearly with $\Omega$ as $G^{\prime\prime}\sim\Omega$. At very low frequencies, this behavior changes to $G^{\prime\prime}\sim\Omega^{\alpha}$ with $\alpha<1$, continuing as $\Omega$ approaches the zero-frequency limit [68].
@@
In the evaluation of viscosity using the NALD approach (Eq. 6), we perform the integration over both real and imaginary branches of $\omega$. As shown in Ref.[59], it is important to include the imaginary modes in the calculation to capture the effect of temperature. A fraction of small-frequency modes below a cutoff frequency $\omega_{\rm min}$ is discarded during the integration, because these modes are unphysical due to the finite system size, as the system cannot support propagating modes corresponding to frequencies below the speed of sound in the medium. Such a minimum frequency can be estimated as $\omega_{\rm min}={2\pi}/{L}\sqrt{{G_{s}}/{\rho}}$, where $L$ is the size of the simulation box, $\rho$ is the mass density, $G_{s}$ is the zero-frequency shear modulus [60]. The zero-frequency shear modulus is calculated by subtracting the shear stress fluctuation from the infinite-frequency shear modulus ($G_{\infty}$), i.e., $G_{s}=G_{\infty}-(V/k_{B}T)\left(\langle\sigma_{xy}^{2}\rangle-\langle\sigma_{xy}\rangle^{2}\right)$, cf. Ref.[81]. The infinite-frequency shear modulus is calculated by deforming the system by a small shear-strain in the linear response regime, and the associated shear-stress is measured including both kinetic and virial components.
@@
Another way to calculate the complex modulus $G^{*}$ is through non-equilibrium molecular dynamics simulations under oscillatory shear [59, 60]. In this method, a small-amplitude oscillatory shear strain $\gamma(t)=\gamma_{0}\sin(\Omega t)$ is applied to the system, and the corresponding stress response $\sigma(t)=\sigma_{0}\sin(\Omega t+\delta)$ is measured, where $\Omega$ is the deformation frequency, and $\delta$ is the phase difference between strain and stress. The $G^{\prime}$ and $G^{\prime\prime}$, which constitute the complex modulus $G^{*}$, are calculated using the relations $G^{\prime}=\left({\sigma_{0}}/{\gamma_{0}}\right)\cos\delta,\quad G^{\prime\prime}=\left({\sigma_{0}}/{\gamma_{0}}\right)\sin\delta$. In these simulations, the strain amplitude $\gamma_{0}$ is kept very small and constant, and a thermostat is applied to maintain a constant temperature during deformation. It has been observed that the data for $G^{\prime}$ and $G^{\prime\prime}$ obtained from the small amplitude oscillatory shear (SAOS) simulations are consistent with the NALD approach at sufficiently higher frequencies. Within the available computational resources, generally, it is not possible to probe the response via SAOS at smaller frequencies accessible in experiments [60]. Therefore, the estimation of viscosity via SAOS, which requires measurement of $G^{\prime\prime}$ at small frequencies, becomes not viable.
@@
Figure 4: (a) Plot of $\sigma_{xy}/\dot{\gamma}$ i.e., the ratio of shear-stress to shear-strain rate, as a function of shear-strain $\gamma$ for different temperatures (marked). (b) Plot of shear viscosity $\eta$ as a function of shear rate $\dot{\gamma}$ at different temperatures (marked).
Figure 5: Comparison of temperature-dependent viscosity obtained using NALD (filled diamonds), Green-Kubo (open circles), and non-equilibrium molecular dynamics (filled squares) methods. The solid line represents the Vogel-Fulcher-Tammann (VFT) fit, confirming the glass transition temperature $T_{g}\simeq 0.4$.
To ensure the convergence of GK integral, we calculate the viscosity as a function of time by performing numerical integration. In Fig. 3(b) we plot viscosity as a function of time at different temperatures. From the figure, we observe that the time-dependent viscosity saturates, making it reliable to extract the viscosity in the long-time plateau region dependent on temperature. At higher temperatures in the liquid state, the saturation region is reached quickly, whereas in the supercooled region with the decrease in temperature, it becomes increasingly difficult to achieve the long-time plateau. We have used these viscosity calculations as a reference to compare the viscosity that we get via the theoretical NALD route.
@@
In the supercooled regime, polymer melts exhibit significant shear thinning [16, 68], characterized by a substantial reduction in viscosity ($\eta$) under high shear rates [82, 83]. To avoid shear thinning effects across the entire temperature range studied, and in Fig. 4(b), we plotted the viscosity as a function of shear strain rate at different temperatures. At higher temperatures, a shear strain rate of $10^{-3}$ is sufficient to obtain a reliable viscosity estimate, where viscosity remains independent of the shear strain rate, thus characterizing the Newtonian regime. However, in the lower temperature regime, shear thinning is observed up to a shear strain rate of $10^{-4}$. In this temperature regime, viscosity depends on the shear rate, exhibiting either shear thinning or shear thickening behavior, which is characteristic of the non-Newtonian regime. The plateau region of viscosity in Fig. 4(b) at different temperatures were extracted and compared with other viscosity data in Fig. 5.
In Fig. 5, the viscosity is plotted as a function of temperature and compared with values obtained from different methods. Open circles represent the results from the Green-Kubo method, filled squares correspond to non-equilibrium MD, and close diamonds are calculated using the NALD method (Eq.6). The NALD method is particularly important for the lower temperature region near the glass transition temperature, where the viscosity exhibits a dramatic increase. The solid line represents the Vogel-Fulcher-Tammann (VFT) fit, given by the equation $\eta=\eta_{0}\exp\left[{A}/{(T-T_{g})}\right]$ where $A=0.75$ is a temperature-independent parameter, $\eta_{0}=15$ represents the microscopic viscosity, and $T_{g}\simeq 0.4$ is the glass transition temperature. The results show a very good agreement in the supercooled region, while some deviations are observed above the melting temperature. However, the overall trend is well captured across the entire temperature range, demonstrating the robustness and consistency of our approach.

### v3

@@ II.2 Green-Kubo method for viscosity calculation
The Green-Kubo (GK) method Green (1954); Kubo (1957) is the most widely used technique to calculate transport coefficients, such as heat conductivity Vaibhav et al. (2020), shear viscosity, and bulk viscosity, relating the time correlation functions to the corresponding fluxes or tensors in thermal equilibrium. To calculate the shear viscosity using this method, at first the autocorrelation function of stress $\sigma_{\alpha\beta}$ (with $\alpha,\beta=x,y,z$) is calculated as a function of time with the equilibrium trajectory of the system. The viscosity is then obtained by integrating the stress autocorrelation function over time, and the expression is given by Helfand (1960)
@@ II.3 Viscosity via non-equilibrium molecular dynamics
Another very popular method to estimate viscosity is via non-equilibrium molecular dynamics (NEMD) Heyes et al. (2018). In this approach, the system is subjected to shear deformation at a rate $\dot{\gamma}$, and the resulting shear stress is measured as the deformation progresses. Initially, the response is linear, followed by a nonlinear regime, until the system reaches a steady state where the shear stress fluctuates around a mean value. If the system is sheared along the $xy$-plane, the stress response is given by $\sigma_{xy}$, and the steady-state average shear stress for a given shear rate is denoted as $\langle\sigma_{xy}\rangle_{\rm S}$. The shear viscosity $\eta$ is then defined as: $\eta=\langle\sigma_{xy}\rangle_{\rm S}/\dot{\gamma}$.
When viscosity is plotted as a function of shear rate, its behavior depends on the nature of the fluid. If viscosity decreases with increasing shear rate, the material exhibits shear-thinning behavior Hara et al. (2025); Heyes et al. (2018), whereas an increase in viscosity with shear rate indicates shear-thickening behavior Morris (2020). A constant viscosity across shear rates characterizes Newtonian behavior. For most of the complex fluids, shear-thinning/shear-thickening vanishes at sufficiently small shear-rates and viscosity becomes a constant independent of shear-rate. Such a viscosity is called zero-shear viscosity given by,
@@ III Results and Discussion
We applied the framework discussed in Section II.1, utilizing the microscopic formula Eq. (6) to establish the relationship between viscosity and normal mode analysis. The vibrational density of states is plotted in Fig. 1(a) for three different temperatures. Since our system is at a finite temperature, the vDOS has a significant number of imaginary modes (conventionally shown on the negative branch of $\omega$) along with the real modesKeyes (1994); Stratt (1995); Zhang et al. (2019). The imaginary modes appear due to the presence of saddles in the energy landscape. As temperature decreases, the number of imaginary modes also decreases, whereas the number of real modes increases. There are two peaks on the positive side of the frequency axis, largest peak occurring at a smaller frequency (close to $\omega=10$) is attributed to the vibration arising from the LJ interactions between beads, while the smaller peak is slightly broader and occurs at higher frequency (close to $\omega=50$), primarily attributed to the vibrations corresponding to non-breakable FENE bonds Jain and de Pablo (2004); Milkus et al. (2018). As expected, $g(\omega)$ exhibits a linear behavior at low frequencies, consistent with both theoretical predictions Zaccone and Baggioli (2021) and experimental observations Stamper et al. (2022); Vaibhav et al. (2025).
@@
Figure 1: (a) Vibrational density of states $g(\omega)$ and (b) affine force field correlator $\Gamma(\omega)$ as a function of normal mode frequency $(\omega)$ at different temperatures (marked). The numerical calculation is done for the KG polymer system described in the text. The quantities are in Lennard-Jones units.
@@
In Fig. 1(b), the affine force field correlator $\Gamma(\omega)$ is plotted as a function of the normal mode frequency $\omega$ for three different temperatures. We observe that $\Gamma(\omega)$ starts increasing rapidly for $\omega>20$ and exhibits a temperature-dependent peak near $\omega\simeq 70$. The peak intensity decreases as the temperature decreases and shifts towards lower frequencies. The decay of the correlator function is also temperature-dependent; faster decay is observed at smaller temperatures. At low frequencies, $\Gamma(\omega)$ is expected to follow the analytical behaviorZaccone and Scossa-Romano (2011); Milkus and Zaccone (2017) $\Gamma\sim\omega^{2}$.
Figure 2: Dependence of the (a) real ($G^{\prime}$, storage modulus) and (b) imaginary ($G^{\prime\prime}$, loss modulus) parts of the complex viscoelastic modulus $(G^{*})$ on external deformation frequency $\Omega$ at different temperatures (marked). Lines represent NALD results and symbols represent MD results obtained from oscillatory shear. The color of the symbols matches that of the lines for the same temperature (only $T=0.6$ and $1.0$ are reported for the MD). The MD simulations are in parameter-free agreement with NALD predictions up to the accessible frequency range. The quantities are in Lennard-Jones units.
@@
Figure 3: (a) Stress autocorrelation function at different temperatures (marked). Inset shows the tail part of the correlation function. (b) Plot of Green-Kubo viscosity as a function of time at different temperatures (marked). The quantities are in Lennard-Jones units.
@@
Using the information of $g(\omega)$ and $\Gamma(\omega)$ we calculate the viscoelastic modulus as a function of external frequency $\Omega$ in the liquid and supercooled regions. In Fig. 2 we plot real and imaginary parts, $G^{\prime}$ and $G^{\prime\prime}$, which represent the storage and loss moduli respectively, as a function of external frequency $\Omega$ at different temperatures. $G^{\prime}$ exhibits specific features: there exists two different resonating frequencies where the response is significantly high with respect to the response at nearby frequencies. The first resonating peak, at a lower frequency, occurs at a value of $\Omega$ that matches with the vibrational eigenfrequency of the LJ interaction. Below this resonance frequency, $G^{\prime}$ decreases monotonically upon decreasing the frequency $\Omega$. The second resonance peak is higher than the previous peak and occurs at a larger value of $\Omega$, close to the characteristic vibration corresponding to FENE bondsPalyulin et al. (2018); Milkus et al. (2018). Since these bonds are non-breakable and more energetic, the mechanical response at $\Omega$ corresponding to FENE bonds is strongest. At the higher-frequency side, greater than any resonating frequencies, there exists a temperature-dependent plateau, representing the high-frequency modulus dominated by purely affine displacements.
@@
In Fig. 2(b), we plot $G^{\prime\prime}$ as a function of $\Omega$ at three temperatures. We observe that the value of $G^{\prime\prime}$ becomes very small at both low and large frequencies. It exhibits a broad peak at $\Omega\sim 70$, and below this frequency $G^{\prime\prime}$ decreases, forming a plateau having a wider width at smaller temperatures due to slower decay in the low-frequency regime. Subsequently, $G^{\prime\prime}$ decreases linearly with $\Omega$ as $G^{\prime\prime}\sim\Omega$ and continues as $\Omega$ approaches the zero-frequency limit Hara et al. (2025).
@@
In the evaluation of viscosity using the NALD approach (Eq. 6), we perform the integration over both real and imaginary branches of $\omega$. As shown in Ref.Palyulin et al. (2018), it is important to include the imaginary modes in the calculation to capture the effect of temperature. A fraction of small-frequency modes below a cutoff frequency $\omega_{\rm min}$ is discarded during the integration, because these modes are unphysical due to the finite system size, as the system cannot support propagating modes corresponding to frequencies below the speed of sound in the medium. Such a minimum frequency can be estimated as $\omega_{\rm min}={2\pi}/{L}\sqrt{{G_{s}}/{\rho}}$, where $L$ is the size of the simulation box, $\rho$ is the mass density, $G_{s}$ is the zero-frequency shear modulus Vaibhav et al. (2024). The zero-frequency shear modulus is calculated by subtracting the shear stress fluctuation from the infinite-frequency shear modulus ($G_{\infty}$), i.e., $G_{s}=G_{\infty}-(V/k_{B}T)\left(\langle\sigma_{xy}^{2}\rangle-\langle\sigma_{xy}\rangle^{2}\right)$, cf. Ref.Fuereder and Ilg (2015). The infinite-frequency shear modulus is calculated by deforming the system by a small shear-strain in the linear response regime, and the associated shear-stress is measured including both kinetic and virial components.
@@
Another way to calculate the complex modulus $G^{*}$ is through non-equilibrium molecular dynamics simulations under oscillatory shear Palyulin et al. (2018); Vaibhav et al. (2024); Sirk et al. (2016). In this method, a small-amplitude oscillatory shear strain $\gamma(t)=\gamma_{0}\sin(\Omega t)$ is applied to the system, and the corresponding stress response $\sigma(t)=\sigma_{0}\sin(\Omega t+\delta)$ is measured, where $\Omega$ is the deformation frequency, and $\delta$ is the phase difference between strain and stress. The $G^{\prime}$ and $G^{\prime\prime}$, which constitute the complex modulus $G^{*}$, are calculated using the relations $G^{\prime}=\left({\sigma_{0}}/{\gamma_{0}}\right)\cos\delta,\quad G^{\prime\prime}=\left({\sigma_{0}}/{\gamma_{0}}\right)\sin\delta$. In these simulations, the strain amplitude $\gamma_{0}$ is kept very small and constant, and a thermostat is applied to maintain a constant temperature during deformation. It has been observed that the data for $G^{\prime}$ and $G^{\prime\prime}$ obtained from the small amplitude oscillatory shear (SAOS) simulations are consistent with the NALD approach at sufficiently higher frequencies. Within the available computational resources, generally, it is not possible to probe the response via SAOS at smaller frequencies accessible in experiments Vaibhav et al. (2024). Therefore, the estimation of viscosity via SAOS, which requires measurement of $G^{\prime\prime}$ at small frequencies, becomes not viable.
We validate the NALD predictions of the viscoelastic modulus against non-equilibrium molecular dynamics (MD) simulations of oscillatory shear, where the $xy$-plane of the simulation box is sheared with an amplitude $\gamma_{0}=0.01$ and frequency $\Omega$ while measuring the stress response $\sigma_{xy}$ as a function of time at temperatures $T=1.0,0.6$. For a reliable estimate, we average the stress response over $10$ to $30$ independent trajectory runs, after excluding the first $30$ cycles of shear deformation. We use stress data and fit trigonometric functions to extract $\sigma_{0}$ and the phase shift $\delta$, allowing us to compute $G^{\prime}$ and $G^{\prime\prime}$. In Figs.  2(a) and 2(b), we compare the storage and loss moduli, as a function of $\Omega$ at temperatures $T=1.0$ and $0.6$, obtained from the NALD approach (lines) with the MD results (symbols). The comparison shows a consistent parameter-free match between NALD and MD down to the accessible frequency range. No adjustable parameter is used in the comparison, because the value of $\nu$ used in the NALD equations and in the Langevin thermostat of the MD simulations is the same.
@@
Figure 4: (a) Plot of $\sigma_{xy}/\dot{\gamma}$ i.e., the ratio of shear-stress to shear-strain rate, as a function of shear-strain $\gamma$ for different temperatures (marked). (b) Plot of shear viscosity $\eta$ as a function of shear rate $\dot{\gamma}$ at different temperatures (marked). The quantities are in Lennard-Jones units.
Figure 5: Comparison of temperature-dependent viscosity obtained using NALD (filled diamonds), Green-Kubo (open circles), and non-equilibrium molecular dynamics (filled squares) methods. The solid line represents the Vogel-Fulcher-Tammann (VFT) fit, confirming the glass transition temperature $T_{g}\simeq 0.4$. The quantities are in Lennard-Jones units.
In Fig. 3(b) we plot viscosity as a function of time at different temperatures. From the figure, we observe that the time-dependent viscosity saturates, making it reliable to extract the viscosity in the long-time plateau region dependent on temperature. At higher temperatures in the liquid state, the saturation region is reached quickly, whereas in the supercooled region with the decrease in temperature, it becomes increasingly difficult to achieve the long-time plateau. We have used these viscosity calculations as a reference to compare the viscosity from the theoretical NALD route.
@@
In the supercooled regime, polymer melts exhibit significant shear thinning Datta et al. (2023); Hara et al. (2025), characterized by a substantial reduction in viscosity ($\eta$) under high shear rates Bonn et al. (2017); Mizuno et al. (2024). To avoid shear thinning effects across the entire temperature range studied, and in Fig. 4(b), we plotted the viscosity as a function of shear strain rate at different temperatures. At higher temperatures, a shear strain rate of $10^{-3}$ is sufficient to obtain a reliable viscosity estimate, where viscosity remains independent of the shear strain rate, thus characterizing the Newtonian regime. However, in the lower temperature regime, shear thinning is observed up to a shear strain rate of $10^{-4}$. In this temperature regime, viscosity depends on the shear rate, exhibiting either shear thinning or shear thickening behavior, which is characteristic of the non-Newtonian regime. The plateau region of viscosity in Fig. 4(b) at different temperatures were extracted and compared with other viscosity data in Fig. 5.
In Fig. 5, the viscosity is plotted as a function of temperature and compared with values obtained from different methods. Open circles represent the results from the Green-Kubo method, filled squares correspond to nonequilibrium MD (NEMD), and close diamonds are calculated using the NALD method (Eq.6). The NALD method is particularly important for the lower temperature region near the glass transition temperature, where the viscosity exhibits a dramatic increase Krausser et al. (2015). The solid line represents the Vogel-Fulcher-Tammann (VFT) fit Mauro et al. (2009), given by the equation $\eta=\eta_{0}\exp\left[{A}/{(T-T_{g})}\right]$ where $A=0.75$ is a temperature-independent parameter, $\eta_{0}=15$, and $T_{g}\simeq 0.4$ is the glass transition temperature. The results show a very good agreement in the supercooled region, while some deviations are observed above the melting temperature. However, the overall trend is well captured across the entire temperature range, demonstrating the robustness and consistency of our approach.